probability question: a class survey carried out in a school showed that 63â„… of the students played sports, 27â„… played musical instruments while 20â„… played neither sports musical instruments. is the event exclusive? Find the probability of a student playing sport or musical instruments or both. Find the probability of playing both musical instrument and sport
1
Expert's answer
2019-11-07T10:11:19-0500

Let "S=" the event "student is playing sport", "M=" the event "student is playing musical instruments"

Two events "A" and "B" are mutually exclusive if they cannot occur at the same time


"P(A\cap B)=0"

Given that "P(S)=0.63, P(M)=0.27, P(\overline{S\cup M})=0.2"

Find the probability of a student playing sport or musical instruments or both


"P(S\cup M)=1-P(\overline{S\cup M})=1-0.2=0.8"

The probability of a student playing sport or musical instruments or both is "0.8"

Find the probability of playing both musical instrument and sport


"P(S\cup M)=P(S)+P(M)-P(S\cap M)"

"P(S\cap M)=P(S)+P(M)-P(S\cup M)"


"P(S\cap M)=0.63+0.27-0.8=0.1"

The probability of a student playing both musical instrument and sport is "0.1"

Hence the events "student is playing sport" and "student is playing musical instruments" are not mutually exclusive.